# Integrable deformations of dimensionally reduced gravity

### v1

@@ D.2 Coupling to $D=2$ dilaton-gravity
00 0 $=\mathrm{d}\star\mathrm{d}\eta-\frac{\beta^{\prime}(\eta)}{\beta(\eta)}\mathrm{d}\eta\wedge\star\mathrm{d}\eta.$
00 0 $=\mathrm{d}\left(\frac{\eta}{\bar{\rho}}\right),\qquad\Rightarrow\qquad\bar{\rho}=C\eta$
The dynamics of $\sigma$ (i.e. the equation of motion from varying w.r.t. to $\rho$), is not reproduced from the Lax connection with spacetime dependent spectral parameter and we neglect it here. This equation of motion has the same for as in the coset case (4.6). As is the undeformed case, one would expect that the equations of motion for $\sigma$ would follow from a combination of the Virasoro constraints and the equations of motion for $\rho$ and the matter part. As we consider this model as coming from a higher dimensional gravity theory, the derivation of $D=2$ Virasoro constraints should be related to this higher-dimensional origin. This higher-dimensional origin is not at all clear for the Yang-Baxter deformation, as discussed in the main text. Hence, the derivation of the Virasoro constraints goes beyond the scope of the present paper, and we do not consider the dynamics of $\sigma$ here.
@@
$\mathrm{d}\rho=\frac{\mathrm{d}\bar{\rho}}{\bar{\beta}(\bar{\rho})}$ (D.15)
In this case $\beta=(1-c^{2}\eta^{2})^{2}$. This can be integrated to:

### v2

@@ D.2 Coupling to $D=2$ dilaton-gravity
00 0 $=\mathrm{d}\star\mathrm{d}\eta-\frac{\beta^{\prime}(\eta)}{\beta(\eta)}\mathrm{d}\eta\wedge\star\mathrm{d}\eta,$
00 0 $=\mathrm{d}\left(\frac{\eta}{\bar{\rho}}\right),\qquad\Rightarrow\qquad\bar{\rho}=C\eta\,.$
The dynamics of $\sigma$ (i.e. the equation of motion from varying w.r.t. to $\rho$), is not reproduced from the Lax connection with spacetime dependent spectral parameter and we neglect it here. This equation of motion has the same form as in the coset case (4.6). As is the undeformed case, one would expect that the equations of motion for $\sigma$ would follow from a combination of the Virasoro constraints and the equations of motion for $\rho$ and the matter part. As we consider this model as coming from a higher dimensional gravity theory, the derivation of $D=2$ Virasoro constraints should be related to this higher-dimensional origin. This higher-dimensional origin is not at all clear for the Yang-Baxter deformation, as discussed in the main text. Hence, the derivation of the Virasoro constraints goes beyond the scope of the present paper, and we do not consider the dynamics of $\sigma$ here.
@@
$\mathrm{d}\rho=\frac{\mathrm{d}\bar{\rho}}{\bar{\beta}(\bar{\rho})}.$ (D.15)
In this case $\beta=(1-c^{2}\bar{\rho}^{2})^{2}$. This can be integrated to: